OpenStudy (anonymous):

Hugh Duffy holds the highest recorded single-season batting average at 0.440. Over his ten year career, his batting average was 0.324. What is the percent of increase of his best season over his career average?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@dockworker

OpenStudy (anonymous):

@brainshot3

OpenStudy (anonymous):

lets call x1 the original value and x2 the new value the formula to find percent change is: \[\frac{x_{2}-x_{1}}{x_{1}}*100\]

hero (hero):

\[\frac{|.440 - .324|}{.440} \times 100 = \frac{.116}{.440} \times 100 = 26.36\]

hero (hero):

Dockworker I posted the formula in previous sections

hero (hero):

sorry, you weren't there

OpenStudy (anonymous):

actually for percent increase, you would use absolute value as hero did. i didn't see that part

OpenStudy (anonymous):

oh, someone just looking for answers?

hero (hero):

I wasn't saying that. I'm not really sure. I think the person wants to understand it. Might be checking answers. You never know.

OpenStudy (anonymous):

Um.... I am kinda lost....

hero (hero):

I can imagine

OpenStudy (anonymous):

just remember if they want the percent increase or decrease, its always positive, if they want percent change, answer can be positive or negative

OpenStudy (anonymous):

well, always positive assuming it actually increases or decreases as they ask

OpenStudy (anonymous):

nevermind, probably more confusing now

hero (hero):

If they say find percent change, then you have to figure out for yourself if the change is an increase or decrease. However, if they say percent increase or decrease, then you don't have to worry about confusing yourself.

OpenStudy (anonymous):

yes, that's a better way to put it.

OpenStudy (anonymous):

so... how do i figure it out